The price of a magazine subscription rose 4% this year. If the subscription now costs $26, how much did it cost last year?
I used $26 divided by .04 = $850. I took $26.00 and subtracted $$8.50 to equal $19.50 which looks about right to me, but the decimal was in the wrong place, so I don't know what formula I should have used, because this one is definitely wrong.
Please tell me the correct formula to use! (The problem below is not a second problem, just another problem with a known answer that I used to attempt to figure out the formula for the actual question above)
I've been using the problem: "The price of a computer fell 20% this year. If the computer now costs $750, how much did it cost last year?" to figure out how to solve the above problem since I know the correct answer to the computer problem is $ 937.50, but I can't solve it.
I used:
$750 divided by 20= $150, then added $750 to $150 to equal $900, which is $37.50 short of the book's answer of $937.50. Each way I've tried brings me to $900.

Answer by richwmiller(17219)

magazine problem
1+.04=1.04
1.04x=26
x=25
another way
.04=1/25
1/25*25=1
add 1+25=26
computer problem
1-.20=.80
.80x=750
x=937.50
These problems used to drive me nuts. The solution is simple once you know how.